pcl5 dissociates 30% at 127degree celsius and 1 atm pressure. Calculate kp and kc of this disocciation reaction.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
PCl5 -------> PCl3 + Cl2
 
1 - 0.3          0.3        0.3
 
Kc = 0.3 *0.3/0.7
Kc = 0.09/0.7 = 0.12857
 
 
Regards
Arun